### v2

# Bifurcation mechanism of quasihalo orbit from Lissajous orbit

Abstract
This paper presents a general analytical method to describe the center manifolds of collinear libration points in the Restricted Three-body Problem (RTBP). It is well-known that these center manifolds include Lissajous orbits, halo orbits, and quasihalo orbits. Previous studies have traditionally treated these orbits separately by iteratively constructing high-order series solutions using the Lindstedt-Poincaré method. Instead of relying on resonance between their frequencies, this study identifies that halo and quasihalo orbits arise due to intricate coupling interactions between in-plane and out-of-plane motions. To characterize this coupling effect, a novel concept, coupling coefficient $\eta$, is introduced in the RTBP, incorporating the coupling term $\eta\Delta x$ in the $z$-direction dynamics equation, where $\Delta$ represents a formal power series concerning the amplitudes. Subsequently, a uniform series solution for these orbits is constructed up to a specified order using the Lindstedt-Poincaré method. For any given paired in-plane and out-of-plane amplitudes, the coupling coefficient $\eta$ is determined by the bifurcation equation $\Delta=0$. When $\eta$ = 0, the proposed solution describes Lissajous orbits around libration points. As $\eta$ transitions from zero to non-zero values, the solution describes quasihalo orbits, which bifurcate from Lissajous orbits. Particularly, halo orbits bifurcate from planar Lyapunov orbits if the out-of-plane amplitude is zero. The proposed method provides a unified framework for understanding these intricate orbital behaviors in the RTBP.
1 Introduction
[…]
The Restricted Three-body Problem (RTBP) serves as a fundamental dynamical model to study the motion of asteroids or satellites under the gravitational influence of two primaries [1, 2, 3]. The RTBP has five equilibrium points called the Lagrange points $L_{i}$ ($i$ = 1, 2, …, 5), among which three so-called collinear libration points are unstable, and the remaining two are triangular libration points. The linear behavior of three collinear libration points is of the type center $\times$ center $\times$ saddle [4, 5, 6]. The center manifold of three collinear libration points includes Lissajous orbits around libration points, halo orbits, and quasihalo orbits around halo orbits [7]. These libration point orbits play a key role in astronomy and many space science missions [8, 9, 10, 11].
Numerical methods (numerical integration, optimization technology) and semi-analytical methods (parameterization method, normal form, Lindstedt-Poincaré method) are often used to study the dynamics in the center manifolds of collinear libration points. In most cases, numerical methods still require accurate analytical initial values. On the other hand, high-order analytical solutions explicitly and accurately describe the dynamics around libration points, which is very useful for the preliminary space mission design. Farquhar initially proposed the third-order solution for small-amplitude Lissajous orbits and introduced the concept of “halo orbit” [12, 13]. Richardson [14] presented a third-order analytical solution of halo orbits around collinear libration points in the RTBP. He employed the Lindstedt-Poincaré method to construct this solution by introducing a correction term to modify the out-of-plane frequency. By means of the Lindstedt-Poincaré method, Jorba and Masdemont [4, 15] obtained high-order series solutions of Lissajous orbits and halo orbits around collinear libration points with the help of ad hoc algebraic manipulators. Celletti and Luo et al. [16, 17] conducted an analytical study of the Lissajous and halo orbits around collinear points in RTBP for arbitrary mass ratios, utilizing a canonical transformation procedure. Paez and Guzzo [18] provided a semi-analytical construction of halo orbits and halo tubes in the elliptic RTBP by implementing a nonlinear Floquet–Birkhoff resonant normal form. However, these analytical methods cannot yield high-order series solutions of quasihalo orbits. Gómez et al. [19] took a halo orbit with a constant amplitude as a backbone (periodic-varying equilibrium point), and further employed Floquet transformation and Lindstedt-Poincaré method to construct quasihalo orbits. This method necessitates specifying the magnitude of the reference halo orbit and using Floquet transformation, and does not directly provide a complete analytical expression for quasihalo orbits.
So far, there is no explicit solution to uniformly describe Lissajous orbits, halo orbits and quasihalo orbits in the center manifolds, although some semi-analytical methods, such as normal form [4] and parameterization methods [6], or numerical methods, such as multiple Poincaré section [20], can provide a comprehensive description for the center manifolds of RTBP. References [14, 4, 15] successfully gave the third or higher-order series solutions of the Lissajous and halo orbits, yet they failed to construct a high-order series solution of the quasihalo orbits. Since the discovery of halo orbits in 1969, the astrodynamics community has traditionally attributed their origin to bifurcations from planar Lyapunov orbits when the in-plane and out-of-plane frequencies are in a 1:1 resonance. In this paper, we also contend that halo orbits indeed arise from bifurcations of planar Lyapunov orbits. However, we employ a new mechanism, coupling-induced bifurcation, to explain the generation of halo and quasihalo orbits. Consider a two-degree-of-freedom nonlinear dynamical system similar to the part of the center manifolds in the RTBP, but in this system, the degrees of freedom are decoupled, i.e., ${\left[{\ddot{x},\ddot{y}}\right]^{\rm{T}}}={\left[{f(x),f(y)}\right]^{\rm{T}}}$. It is evident that bifurcations of periodic orbits do not occur even under frequency resonance. Drawing a distinction between this hypothetical dynamical system and the RTBP, it can be inferred that the true trigger for the emergence of halo orbits through planar Lyapunov orbit bifurcation is the nonlinear coupling effect between the two degrees of freedom, rather than a resonance between the two frequencies. Considering the nonlinear coupling of in-plane and out-of-plane motions, this paper introduces the concept of a coupling coefficient $\eta$ into the RTBP for the first time. Additionally, a new correction term $\eta\Delta x$ is incorporated into the RTBP equation to construct a unified semi-analytical solution for the center manifolds of collinear libration points in the RTBP. Here, $\Delta$ is a power series concerning in-plane and out-of-plane amplitudes. When the coupling coefficient $\eta$ is zero, the correction item is inactive, and the series solution describes Lissajous orbits around libration points. Otherwise, the series solution describes quasihalo orbits, of course, including halo orbits. In this way, the center manifolds of collinear libration points in the RTBP are uniformly described with a high-order series solution.
The contribution of this paper lies in the novel proposition of the coupling-induced bifurcation mechanism to explain the generation of quasihalo orbits. This marks the first complete approximate analytical solution for quasihalo orbits, providing a comprehensive analytical solution for the entire central manifold, including Lissajous orbits, halo orbits, and quasihalo orbits. Additionally, the practical convergence domain of the Lindstedt-Poincaré expansion of quasihalo orbits is presented. Lastly, the proposed method transforms the dynamic orbit bifurcation problem of vector fields into a static bifurcation problem for the solutions of the bifurcation equation. This method is applicable for general analytical bifurcation analysis in dynamical systems.
The remainder of this paper is organized as follows. Section 2 introduces the dynamical model of the RTBP. In section 3, a general analytical construction method for the center manifolds of collinear libration points in the RTBP is presented. Results and discussions are provided in Section 4. Finally, Section 5 makes some concluding remarks.
2 Dynamical model
This section introduces the dynamical model of the classical RTBP. This model serves as a good approximation for the motion of an infinitesimal particle (asteroids or spacecraft) under the gravitational attraction of two primaries. The attraction of the particle on the primaries is neglected so the two primaries rotating around their common center of mass in a Kepler orbit. Here, our focus lies on the circular orbit, i.e., a Kepler orbit with zero eccentricity. The motion of the particle is usually described in a synodic coordinate system. In this system, the origin is positioned at the centroid of the two primaries. The $X$-axis points from the smaller primary to larger primary, the $Z$-axis is perpendicular to the plane of the Kepler orbit and positive in the direction of the angular momentum, and the $Y$-axis completes a right-hand triad. Consequently, the position vector of the smaller primary and larger primary is ($-\mu$, 0, 0) and ($1-\mu$, 0, 0), respectively. Then, the differential equation governing the motion of the particle in the synodic coordinate system is expressed as follows [21, 7]:
$\ddot{X}-2\dot{Y}$ $=\frac{{\partial\Omega}}{{\partial X}}$ (1)
$\ddot{Y}+2\dot{X}$ $=\frac{{\partial\Omega}}{{\partial Y}}$
$\ddot{Z}$ $=\frac{{\partial\Omega}}{{\partial Z}}$
with
$\Omega\left({X,Y,Z}\right)=\frac{1}{2}\left({{X^{2}}+{Y^{2}}}\right)+\frac{{1-\mu}}{{{r_{1}}}}+\frac{\mu}{{{r_{2}}}}+\frac{1}{2}\mu\left({1-\mu}\right)$ (2)
where $\mu=m_{2}/(m_{1}+m_{2})$ is the mass parameter of the system. $m_{1}$ and $m_{2}$ are the mass of the smaller primary and larger primary, respectively. $r_{1}$ and $r_{2}$ are the distance from the particle to the smaller primary and larger primary, respectively.
$\begin{array}[]{l}{r_{1}}^{2}={\left({X+\mu}\right)^{2}}+{Y^{2}}+{Z^{2}}\\ {r_{2}}^{2}={\left({X-1+\mu}\right)^{2}}+{Y^{2}}+{Z^{2}}\end{array}$ (3)
This model has a Jacobi integral
$C=2\Omega-({\dot{X}^{2}}+{\dot{Y}^{2}}+{\dot{Z}^{2}})$ (4)
As mentioned in the Introduction section, the RTBP has five equilibrium points. In this paper, we focus on the three collinear libration points $L_{1}$, $L_{2}$, and $L_{3}$. Let $\gamma_{i}(i=1,2,3)$ denote the distance from $L_{i}$ to the closet primary. This distance is determined by the unique positive root of the Euler quantic equation [7],
$\begin{array}[]{l}\gamma_{i}^{5}\mp(3-\mu)\gamma_{i}^{4}+(3-2\mu)\gamma_{i}^{3}-\mu\gamma_{i}^{2}\pm 2\mu{\gamma_{i}}-\mu=0,{\rm{}}\qquad\qquad\qquad\quad i=1,2\\ \gamma_{i}^{5}+(2+\mu)\gamma_{i}^{4}+(1+2\mu)\gamma_{i}^{3}-(1-\mu)\gamma_{i}^{2}-2(1-\mu){\gamma_{i}}-(1-\mu)=0,\ i=3\end{array}$ (5)
When focusing on the selected libration point $L_{i}$, a coordinate transformation is performed to move the origin of coordinate system to the libration point $L_{i}$,
$\begin{array}[]{l}X=-{\gamma_{i}}x+\mu-1+{\gamma_{i}},Y=-{\gamma_{i}}y,Z={\gamma_{i}}z,{\rm{}}\ i=1,2\\ X={\gamma_{i}}x+\mu+{\gamma_{i}},Y={\gamma_{i}}y,Z={\gamma_{i}}z,\quad\quad\quad{\rm{}}i=3\end{array}$ (6)
Then, the dynamical model (1) of the RTBP can be reformulated as
$\ddot{x}-2\dot{y}$ $=\frac{{\partial\Omega}}{{\partial x}}$ (7)
$\ddot{y}+2\dot{x}$ $=\frac{{\partial\Omega}}{{\partial y}}$
$\ddot{z}$ $=\frac{{\partial\Omega}}{{\partial z}}$
where
$\Omega\left({x,y,z}\right)$ $=\frac{1}{2}\left({{{\left({\mu-1\mp\gamma\left({x-1}\right)}\right)}^{2}}+{\gamma^{2}}{y^{2}}}\right)$ (8)
$+\frac{{1-\mu}}{{{r_{1}}}}+\frac{\mu}{{{r_{1}}}}+\frac{1}{2}\mu\left({1-\mu}\right)$
In order to construct a high-order series solution for the center manifolds in the RTBP in the following section, the motion equation (7) is expanded in power series using the Legendre polynomials [4, 22],
$\ddot{x}-2\dot{y}-(1+2{c_{2}})x$ $=\frac{\partial}{{\partial x}}\sum\limits_{n\geq 3}{{c_{n}}{\rho^{n}}{P_{n}}\left({\frac{x}{\rho}}\right)}$ (9)
$\ddot{y}+2\dot{x}+({c_{2}}-1)y$ $=\frac{\partial}{{\partial y}}\sum\limits_{n\geq 3}{{c_{n}}{\rho^{n}}{P_{n}}\left({\frac{y}{\rho}}\right)}$
$\ddot{z}+{c_{2}}z$ $=\frac{\partial}{{\partial z}}\sum\limits_{n\geq 3}{{c_{n}}{\rho^{n}}{P_{n}}\left({\frac{z}{\rho}}\right)}$
where $\rho^{2}=x^{2}+y^{2}+z^{2}$, $P_{n}$ is Legendre polynomials, and the constant coefficients $c_{n}(\mu)$ depend on the system parameters $\mu$,
$\begin{array}[]{l}{c_{n}}(\mu)=\frac{1}{{\gamma_{i}^{3}}}\left({{{(\pm 1)}^{n}}+{{(-1)}^{n}}\frac{{(1-\mu)\gamma_{i}^{n+1}}}{{{{(1\mp\gamma)}^{n+1}}}}}\right),\ {\rm{for}}\ {L_{i}},\ i=1,2\\ {c_{n}}(\mu)=\frac{{{{(-1)}^{n}}}}{{\gamma_{i}^{3}}}\left({1-\mu+\frac{{\mu\gamma_{i}^{n+1}}}{{{{(1+\gamma)}^{n+1}}}}}\right),\quad\quad{\rm{for}}\ {L_{i}},\ i=3.\end{array}$ (10)
The left-hand linear terms form the foundational components of the series solution, while the right-hand nonlinear ones are crucial for constructing halo and quasihalo orbits. They are well-known as bifurcations from large-amplitude planar Lyapunov orbit.
3 Analytical construction of center manifolds in RTBP
In this section, the coupling coefficient is precisely defined. A coupling correction term is introduced to the original dynamical equation, which leads to a modified linear solution that accounts for coupling effects. Based on the new linear solution, the Lindstedt–Poincaré method is applied to construct a unified analytical solution for the center manifold near collinear libration points in the RTBP.
3.1 Form of the analytical solution
First, the first-order solution for the center manifolds of collinear libration points in the RTBP should be found. It is well known that the center manifolds include Lissajous orbits, halo orbits, and quasihalo orbits around halo orbits. For Lissajous orbits, the first-order solution naturally arises by solving the linear part of (9),
$\ddot{x}-2\dot{y}-(1+2{c_{2}})x$ $=0$ (11)
$\ddot{y}+2\dot{x}+({c_{2}}-1)y$ $=0$
$\ddot{z}+{c_{2}}z$ $=0.$
The solution of (11) is
$\begin{array}[]{l}x(t)={\rm{}}\alpha\cos({\omega_{0}}t+{\varphi_{1}})\\ y(t)=\kappa\alpha\sin({\omega_{0}}t+{\varphi_{1}})\\ z(t)=\beta\cos({v_{0}}t+{\varphi_{2}})\end{array}$ (12)
where
${\omega_{0}}=\sqrt{\frac{{2-{c_{2}}+\sqrt{9c_{2}^{2}-8{c_{2}}}}}{2}},{v_{0}}=\sqrt{{c_{2}}},\kappa=-\frac{{{\omega_{0}}^{2}+1+{2c_{2}}}}{{2{\omega_{0}}}}$ (13)
$\alpha$ and $\beta$ represent the in-plane and out-of-plane amplitudes, respectively. $\phi_{1}$ and $\phi_{2}$ denote the corresponding phases. As stated in the Introduction section, halo orbits bifurcate from the planar Lyapunov periodic orbits without requiring the in-plane and out-of-plane frequencies to be equal. This phenomenon is primarily due to the nonlinear coupling of in-plane and out-of-plane motions in RTBP. Considering the coupling between in-plane and out-of-plane motions, a natural concept of coupling coefficient $\eta$ can be defined to characterize the degree of coupling between the linear motion in plane and out of plane. Consequently, the basic first-order solution can be reformulated as
$\begin{array}[]{l}x(t)={\rm{}}\alpha\cos({\omega_{0}}t+{\varphi_{1}})\\ y(t)=\kappa\alpha\sin({\omega_{0}}t+{\varphi_{1}})\\ z(t)=\eta\alpha\cos({\omega_{0}}t+{\varphi_{1}})+\beta\cos({v_{0}}t+{\varphi_{2}})\end{array}$ (14)
To derive the solution (14) for (11), it is necessary to introduce a correction term and rewrite it as
$\ddot{x}-2\dot{y}-(1+2{c_{2}})x$ $=0$ (15)
$\ddot{y}+2\dot{x}+({c_{2}}-1)y$ $=0$
$\ddot{z}+{c_{2}}z$ $=\eta{d_{00}}x$
where $d_{00}=c_{2}-{\omega_{0}}^{2}\neq 0$ is a constant correction factor and $\eta d_{00}=0$. It is evident that in the linear case, $\eta d_{00}=0$ only has a unique trivial solution, i.e., $\eta=0$, indicating the absence of coupling between the in-plane and out-of-plane motions. Halo/quasihalo orbits only appears when the coupling coefficient $\eta$ is not-zero in higher-order solutions. Therefore, we extend (15) to higher-order case by introducing a higher-order correction term to the third equation of (9) as follows:
$\ddot{x}-2\dot{y}-(1+2{c_{2}})x$ $=\frac{\partial}{{\partial x}}\sum\limits_{n\geq 3}{{c_{n}}{\rho^{n}}{P_{n}}\left({\frac{x}{\rho}}\right)}$ (16)
$\ddot{y}+2\dot{x}+({c_{2}}-1)y$ $=\frac{\partial}{{\partial y}}\sum\limits_{n\geq 3}{{c_{n}}{\rho^{n}}{P_{n}}\left({\frac{y}{\rho}}\right)}$
$\ddot{z}+{c_{2}}z$ $=\frac{\partial}{{\partial z}}\sum\limits_{n\geq 3}{{c_{n}}{\rho^{n}}{P_{n}}\left({\frac{z}{\rho}}\right)}+\eta\Delta x$
where the higher-order term $\eta\Delta x$ is the product of the coupling coefficient $\eta$, the nonlinear correction factor $\Delta$, and $x$, satisfying $\Delta=\sum\limits{{d_{ij}}{\alpha^{i}}{\beta^{j}}}$ and $\eta\Delta=0$. Since the coefficients $d_{ij}$ of $\Delta$ also include $\eta$, $\Delta$ is a polynomial function of $\eta$. During the process, non-trivial solutions of $\eta$ are determined from the condition $\Delta=0$. When the nonlinear terms in (16) are taken into account, the semi-analytical solution is formulated as formal expansions in powers of $\alpha$ and $\beta$,
$\begin{array}[]{l}x\left(t\right)=\sum\limits_{i+j\geq 1}{\left({\sum\limits_{\left|k\right|\leq i,\left|m\right|\leq j}{{x_{ijkm}}\cos\left({k{\theta_{1}}+m{\theta_{2}}}\right)}}\right){\alpha^{i}}{\beta^{j}}}\\ y\left(t\right)=\sum\limits_{i+j\geq 1}{\left({\sum\limits_{\left|k\right|\leq i,\left|m\right|\leq j}{{y_{ijkm}}\sin\left({k{\theta_{1}}+m{\theta_{2}}}\right)}}\right){\alpha^{i}}{\beta^{j}}}\\ z\left(t\right)=\sum\limits_{i+j\geq 1}{\left({\sum\limits_{\left|k\right|\leq i,\left|m\right|\leq j}{{z_{ijkm}}\cos\left({k{\theta_{1}}+m{\theta_{2}}}\right)}}\right){\alpha^{i}}{\beta^{j}}}\end{array}$ (17)
where ${\theta_{1}}{\rm{=}}{\omega}t+{\varphi_{1}}$,${\theta_{2}}{\rm{=}}\nu t+{\varphi_{2}}$. Considering the presence of high-order terms, the frequencies should also be expanded in the power series of $\alpha$ and $\beta$,
$\omega=\sum\limits_{i,j\geq 0}{{\omega_{ij}}}{\alpha^{i}}{\beta^{j}},v=\sum\limits_{i,j\geq 0}{{v_{ij}}}{\alpha^{i}}{\beta^{j}}$ (18)
Moreover, we have the constraint
$\eta\Delta=\eta\sum\limits_{i+j=n}{{d_{ij}}{\alpha^{i}}{\beta^{j}}}=0$ (19)
which provides the implicit relationship between $\eta$ and $\alpha$ and $\beta$, i.e., $\eta=\eta(\alpha,\beta)$. Equations (17), (18), and (19) together constitute a comprehensive analytical description of the central manifolds around the collinear libration in the RTBP. The process for solving the coefficients $x_{ijkm}$, $y_{ijkm}$, $z_{ijkm}$, $\omega_{ij}$, $v_{ij}$ and $d_{ij}$ is provided in Appendix 3.2, and it is similar to the approach in [4], except for the parts related to the coupling coefficient $\eta$.
When $\eta=0$, the solution (17) represents Lissajous orbits. Specifically, if $\beta=0$, it corresponds to planar Lyapunov orbits; if $\alpha=0$, it corresponds to vertical Lyapunov orbits. When $\eta\neq 0$, the solution (17) represents quasihalo orbits. Particularly, if $\eta>0$ $(\eta<0)$ and $\beta=0$, it results in north (south) halo orbits. No solutions exist if $\alpha=0$, which will be demonstrated in the next section. In summary, the high-order series solution (17) uniformly describes the center manifolds of collinear libration points in the RTBP.
Although the analysis above only considered the coupling of $x$ applied to $z$, different forms of coupling will actually bifurcate into different families of orbits. When the $x$ coupling effect is applied to $z$, the solution of $\Delta=0$ indicates that this family of orbits bifurcates from planar Lyapunov orbits (and the corresponding Lissajous orbit), resulting in the so-called halo and quasihalo orbits. It is known that halo orbits cannot bifurcate from vertical Lyapunov orbits in the RTBP. Therefore, if we consider coupling $z$ to $x$, the equation $\Delta=0$ has no solution. Furthermore, if we consider coupling $\eta\Delta y$ to $z$ or $\eta\Delta z$ to $y$, another family of orbits, known as axis orbit and quasi-axis orbits, will be obtained. It is well-known that axis orbits can be regarded as bifurcating either from planar Lyapunov orbits or from vertical Lyapunov orbits [23].
3.2 Solving for undetermined coefficients
This part provides the process for solving the coefficients $x_{ijkm}$, $y_{ijkm}$, $z_{ijkm}$, $\omega_{ij}$, $v_{ij}$ and $d_{ij}$. From (17) and (18) and (19), it is known that obtaining the analytical expressions for the center manifold near the $L_{1}$ point is contingent upon determining the coefficients $x_{ijkm}$, $y_{ijkm}$, $z_{ijkm}$, $\omega_{ij}$, $v_{ij}$ and $d_{ij}$. Utilizing the Lindstedt–Poincaré method with iterations starting from the linear solution allows for the computation of these coefficients. Now, if the coefficients and the corresponding solution up to order $n-1$ are established, substituting this solution into the right-hand side of (16) yields the three known series up to order $n$, denoted by $p$, $q$, and $r$. The terms up to the $n$-order from these series, $p_{ijkm}$, $q_{ijkm}$, and $r_{ijkm}$ ($i+j=n$), are selected for solving $n$-order unknown coefficients.
Then, the known and unknown terms on the left side of the equation are classified. The first and second derivatives of x are expressed as:
$\begin{array}[]{l}\dot{x}=\omega\frac{{\partial x}}{{\partial{\theta_{1}}}}+\nu\frac{{\partial x}}{{\partial{\theta_{2}}}}\\ \ddot{x}={\omega^{2}}\frac{{{\partial^{2}}x}}{{\partial\theta_{1}^{2}}}+2\omega v\frac{{{\partial^{2}}x}}{{\partial{\theta_{1}}\partial{\theta_{2}}}}+{v^{2}}\frac{{{\partial^{2}}x}}{{\partial\theta_{2}^{2}}}\end{array}$ (20)
Analogous operations can be carried out for $y$ and $z$. For the first-order derivatives and the coupling correction term $\Delta x$, their known and unknown terms can be classified as shown in Table 1, where $i_{f}$ represent the power of frequency terms or coupling term $\Delta$, and $i_{g}$ represents the power of coordinate terms, $\delta_{ij}$ denotes Kronecker function. A similar classification for the second-order derivatives is provided in Table 2.
Table 1: Classification for the derivatives of $x$ and $y$ and coupling correction term $\Delta x$.
$f$ $g$ $\omega(\partial x/\partial{\theta_{1}})$ $v(\partial x/\partial{\theta_{2}})$ $\omega(\partial y/\partial{\theta_{1}})$ $v(\partial y/\partial{\theta_{2}})$ $\Delta x$
Unknown term 0 n $-\omega_{00}kx_{ijkm}$ $-v_{00}mx_{ijkm}$ $\omega_{00}ky_{ijkm}$ $v_{00}my_{ijkm}$ $d_{00}x_{ijkm}$
$n-1$ 1 $-\omega_{i-1j}k\delta_{1k0m}$ 0 $\kappa\omega_{i-1j}k\delta_{1k0m}$ 0 $d_{i-1j}\delta_{1k0m}$
Known term 1,2,…, $n-2$ $n-(ij)_{f}$ $-\omega_{(ij)_{f}}kx_{(ij)_{g}km}$ $-v{(ij)_{f}}mx_{(ij)_{g}km}$ $\omega_{(ij)_{f}}ky_{(ij)_{g}km}$ $v{(ij)_{f}}my_{(ij)_{g}km}$ ${d_{{{(ij)}_{f}}}}{x_{{{(ij)}_{g}}km}}$
Table 2: Classification for the second derivatives of $x$, $y$, and $z$.
$f$ $g$ $\omega^{2}(\partial^{2}[x,y,z]/\partial\theta_{1}^{2})$ $2\omega v(\partial^{2}[x,y,z]/\partial\theta_{1}\theta_{2})$ $v^{2}(\partial^{2}[x,y,z]/\partial\theta_{1}^{2})$
Unknown term 0 $n$ $-\omega_{00}^{2}k^{2}[x,y,z]_{ijkm}$ $-2\omega_{00}v_{00}km[x,y,z]_{ijkm}$ $-v_{00}^{2}m^{2}[x,y,z]_{ijkm}$
$n-1$ 1 $\begin{array}[]{c}-\Omega_{i-1j}\delta_{1k}\delta_{0m}\\ -\Omega_{i-1j}\kappa\delta_{1k}\delta_{0m}\\ -\Omega_{i-1j}\eta\delta_{1k}\delta_{0m}\end{array}$ 0 $\begin{array}[]{c}0\\ 0\\ -N_{(ij-1)}\delta_{0k}\delta_{1m}\end{array}$
Known term 1,2,…, $n-2$ $n-(ij)_{f}$ $-\Omega_{(ij)_{f}}k^{2}[x,y,z]_{(ij)_{g}km}$ $-2\Omega_{(ij)_{f_{1}}}N_{(ij)_{f_{2}}}km[x,y,z]_{(ij)_{g}km}$ $-N{(ij)_{f}}m^{2}[x,y,z]_{(ij)_{g}km}$
Then, all the known terms from Table 1 and Table 2 are combined into the known series on the right-hand side of (16) and they are denoted as ${\bar{p}_{ijkm}}$, ${\bar{q}_{ijkm}}$ and ${\bar{r}_{ijkm}}$. Additionally, for the unknown terms $\Omega_{i-1j}$ and ${\rm N}_{ij-1}$, they also include some known terms, i.e., $\Omega_{i-1j}=2\omega_{00}\omega_{i-1j}+\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}$, ${\rm N}_{ij-1}=2v_{00}v_{ij-1}+\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{v_{{i_{1}}{j_{1}}}}{v_{{i_{2}}{j_{2}}}}}$, where $i_{1}+j_{1}+i_{2}+j_{2}=n-1$ and $C_{n}^{k}$ denotes a combination. These known terms need to be merged into the right-hand side of the equation as well. In conclusion, the linear system of equations for each term of order $n$ is as follows:
$\begin{array}[]{l}-(\varpi_{km}^{2}+1+2{c_{2}}){x_{ijkm}}-2{\varpi_{km}}{y_{xjkm}}-2\left({{\omega_{00}}+\kappa}\right){\omega_{i-1j}}{\delta_{1k}}{\delta_{0m}}={{\bar{\bar{p}}}_{ijkm}}\\ -2{\varpi_{km}}{x_{ijkm}}+({c_{2}}-1-\varpi_{km}^{2}){y_{xjkm}}-2\left({\kappa{\omega_{00}}+1}\right){\omega_{i-1j}}{\delta_{1k}}{\delta_{0m}}={{\bar{\bar{q}}}_{ijkm}}\\ ({c_{2}}-\varpi_{km}^{2}){z_{ijkm}}-{d_{00}\eta}{x_{ijkm}}-2{v_{00}}{v_{ij-1}}{\delta_{0k}}{\delta_{1m}}-2\eta{\omega_{00}}{\omega_{i-1j}}{\delta_{1k}}{\delta_{0m}}-\eta{d_{i-1j}}{\delta_{1k}}{\delta_{0m}}={{\bar{\bar{r}}}_{ijkm}}\end{array}$ (21)
where ${\varpi_{km}}=k{\omega_{00}}+m{v_{00}}$, and
${{\bar{\bar{p}}}_{ijkm}}$ $={{\bar{p}}_{ijkm}}+{\delta_{k1}}{\delta_{0m}}\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}$ (22)
${{\bar{\bar{q}}}_{ijkm}}$ $={{\bar{q}}_{ijkm}}+\kappa{\delta_{k1}}{\delta_{0m}}\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}$
${{\bar{\bar{r}}}_{ijkm}}$ $={{\bar{r}}_{ijkm}}+\eta{\delta_{1k}}{\delta_{0m}}\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}$
$+{\delta_{0k}}{\delta_{1m}}\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{v_{{i_{1}}{j_{1}}}}{v_{{i_{2}}{j_{2}}}}}.$
When $(k,m)\neq(1,0),(k,m)\neq(0,1)$, (21) becomes the regular linear equations (23). The coefficients $x_{ijkm}$, $y_{ijkm}$, and $z_{ijkm}$ can be solved immediately.
$-(\varpi_{km}^{2}+1+2{c_{2}}){x_{ijkm}}-2{\varpi_{km}}{y_{xjkm}}$ $={{\bar{\bar{p}}}_{ijkm}}$ (23)
$-2{\varpi_{km}}{x_{ijkm}}+({c_{2}}-1-\varpi_{km}^{2}){y_{xjkm}}$ $={{\bar{\bar{q}}}_{ijkm}}$
$({c_{2}}-\varpi_{km}^{2}){z_{ijkm}}-{d_{00}\eta}{x_{ijkm}}$ $={{\bar{\bar{r}}}_{ijkm}}.$
When $(k,m)=(1,0)$, $x_{ijkm}$, $y_{ijkm}$, and $z_{ijkm}$ are coupled. Thus, $x_{ijkm}$ can be set zero, $z_{ijkm}=\eta x_{ijkm}=0$. In this case, (21) is turned to linear equations (24). Then, $y_{ijkm}$ and $\omega_{i-1j}$ are solved by the first two equations of (24). $d_{i-1j}$ is obtained from $\eta{d_{i-1j}}=-{\bar{r}_{ijkm}}-\eta\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}$.
$-2{\omega_{00}}{y_{ijkm}}-2\left({{\omega_{00}}+\kappa}\right){\omega_{i-1j}}$ $={{\bar{p}}_{ijkm}}+\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}$ (24)
$+({c_{2}}-1-\omega_{00}^{2}){y_{ijkm}}-2\left({\kappa{\omega_{00}}+1}\right){\omega_{i-1j}}$ $={{\bar{q}}_{ijkm}}+\kappa\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}$
$-\eta{d_{i-1j}}$ $={{\bar{r}}_{ijkm}}{\rm{+}}\eta\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}.$
When $(k,m)=(0,1)$, the coefficient of $z_{ijkm}$ is zero, and therefore $z_{ijkm}$ is set to zero. In this case, (21) simplifies into linear equations (25). Then, $x_{ijkm}$ and $y_{ijkm}$ are solved by the first two equations of (25). $v_{ij-1}$ is obtained from $-2{v_{00}}{v_{ij-1}}={\bar{r}_{ijkm}}+{d_{00}}{x_{ijpq}}+\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{v_{{i_{1}}{j_{1}}}}{v_{{i_{2}}{j_{2}}}}}$.
$-(v_{00}^{2}+1+2{c_{2}}){x_{ijkm}}-2{v_{00}}{y_{ijkm}}$ $={{\bar{p}}_{ijkm}}$ (25)
$-2{v_{00}}{x_{ijkm}}+({c_{2}}-1-v_{00}^{2}){y_{ijkm}}$ $={{\bar{q}}_{ijkm}}$
$-{d_{00}}{x_{ijkm}}-2{v_{00}}{v_{ij-1}}$ $={{\bar{r}}_{ijkm}}+\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{v_{{i_{1}}{j_{1}}}}{v_{{i_{2}}{j_{2}}}}}.$
4 Results
In this section, the third-order analytical solution for central manifolds around the collinear libration points in the RTBP with arbitrary system parameter $\mu$ is derived. Moreover, the construction of the series solution up to a certain order $n$ is implemented for the given system parameter $\mu$, such as the Sun-Earth (S-E) system ($\mu=3.040423398444176\times 10^{-6}$) or Earth-Moon (E-M) system ($\mu=1.2150581918706896\times 10^{-2}$), utilizing the C++ 17 programming language.
4.1 Third-order analytical solution
It is well-known that halo/quasihalo orbits in the RTBP first appear in the third-order series solution. To achieve a comprehensive description of the center manifolds, their third-order analytical solution is derived using the analytical construction method described in Section 3, as follows:
$x$ $=\alpha\cos{\theta_{1}}+\left({{a_{21}}+{a_{22}}{\eta^{2}}}\right){\alpha^{2}}+\left({{a_{23}}+{a_{24}}{\eta^{2}}}\right){\alpha^{2}}\cos 2{\theta_{1}}$ (26)
$+{a_{25}}\eta\alpha\beta\cos({\theta_{1}}+{\theta_{2}})+{a_{26}}\eta\alpha\beta\cos({\theta_{1}}-{\theta_{2}})$
$+{a_{27}}{\beta^{2}}+{a_{28}}{\beta^{2}}\cos 2{\theta_{2}}$
$+({{a_{31}}{\eta^{4}}+{a_{32}}{\eta^{2}}+{a_{33}}}){\alpha^{3}}\cos 3{\theta_{1}}+({{a_{34}}{\eta^{3}}+{a_{35}}\eta}){\alpha^{2}}\beta\cos{\theta_{2}}$
$+({{a_{36}}{\eta^{3}}+{a_{37}}\eta}){\alpha^{2}}\beta\cos(2{\theta_{1}}+{\theta_{2}})$
$+({{a_{38}}{\eta^{3}}+{a_{39}}\eta}){\alpha^{2}}\beta\cos(2{\theta_{1}}-{\theta_{2}})$
$+({{a_{310}}{\eta^{2}}+{a_{311}}})\alpha{\beta^{2}}\cos({\theta_{1}}+2{\theta_{2}})$
$+({{a_{312}}{\eta^{2}}+{a_{313}}})\alpha{\beta^{2}}\cos({\theta_{1}}-2{\theta_{2}})$
$+{a_{314}}\eta{\beta^{3}}\cos{\theta_{2}}+{a_{315}}\eta{\beta^{3}}\cos 3{\theta_{2}},$
$y$ $=\kappa\alpha\sin{\theta_{1}}+({{b_{21}}+{b_{22}}{\eta^{2}}}){\alpha^{2}}\sin 2{\theta_{1}}+{b_{23}}\eta\alpha\beta\sin({\theta_{1}}+{\theta_{2}})$
$+{b_{24}}\eta\alpha\beta\sin({\theta_{1}}-{\theta_{2}})+{b_{25}}{\beta^{2}}\sin 2{\theta_{2}}$
$+({{b_{31}}{\eta^{4}}+{b_{32}}{\eta^{2}}+{b_{33}}}){\alpha^{3}}\sin{\theta_{1}}+({{b_{34}}{\eta^{4}}+{b_{35}}{\eta^{2}}+{b_{36}}}){\alpha^{3}}\sin 3{\theta_{1}}$
$+({{b_{37}}{\eta^{3}}+{b_{38}}\eta}){\alpha^{2}}\beta\sin{\theta_{2}}+({{b_{39}}{\eta^{3}}+{b_{310}}\eta}){\alpha^{2}}\beta sin(2{\theta_{1}}+{\theta_{2}})$
$+({{b_{311}}{\eta^{3}}+{b_{312}}\eta}){\alpha^{2}}\beta sin(2{\theta_{1}}-{\theta_{2}})+({{b_{313}}{\eta^{2}}+{b_{314}}})\alpha{\beta^{2}}\sin{\theta_{1}}$
$+({{b_{315}}{\eta^{2}}+{b_{316}}})\alpha{\beta^{2}}\sin({\theta_{1}}+2{\theta_{2}})$
$+({{b_{317}}{\eta^{2}}+{b_{318}}})\alpha{\beta^{2}}\sin({\theta_{1}}-2{\theta_{2}})$
$+{b_{319}}\eta{\beta^{3}}\sin{\theta_{2}}+{b_{320}}\eta{\beta^{3}}\sin 3{\theta_{2}},$
$z$ $=\eta\alpha\cos{\theta_{1}}+\left({{d_{21}}+{d_{22}}{\eta^{2}}}\right)\eta{\alpha^{2}}+\left({{d_{23}}+{d_{24}}{\eta^{2}}}\right)\eta{\alpha^{2}}\cos 2{\theta_{1}}$
$+({{d_{25}}+{d_{26}}{\eta^{2}}})\alpha\beta\cos({\theta_{1}}+{\theta_{2}})+({{d_{27}}+{d_{28}}{\eta^{2}}})\alpha\beta\cos({\theta_{1}}-{\theta_{2}})$
$+{d_{29}}\eta{\beta^{2}}+{d_{210}}\eta{\beta^{2}}\cos 2{\theta_{2}}+({{d_{31}}{\eta^{5}}+{d_{32}}{\eta^{3}}+{d_{33}}\eta}){\alpha^{3}}\cos 3{\theta_{1}}$
$+({{d_{34}}{\eta^{4}}+{d_{35}}{\eta^{2}}+{d_{36}}}){\alpha^{2}}\beta\cos(2{\theta_{1}}+{\theta_{2}})$
$+({{d_{37}}{\eta^{4}}+{d_{38}}{\eta^{2}}+{d_{39}}}){\alpha^{2}}\beta\cos(2{\theta_{1}}-{\theta_{2}})$
$+\left({{d_{310}}{\eta^{3}}+{d_{311}}\eta}\right)\alpha{\beta^{2}}\cos({\theta_{1}}+2{\theta_{2}})$
$+({{d_{312}}{\eta^{3}}+{d_{313}}\eta})\alpha{\beta^{2}}\cos({\theta_{1}}-2{\theta_{2}})$
$+({{d_{314}}{\eta^{2}}+{d_{315}}}){\beta^{3}}\cos 3{\theta_{2}},$
with the frequencies
$\omega$ $={\omega_{0}}+{\omega_{20}}+{\omega_{02}}={e_{31}}{\alpha^{2}}{\eta^{4}}+\left({{e_{32}}{\alpha^{2}}+{e_{33}}{\beta^{2}}}\right){\eta^{2}}+{e_{34}}{\alpha^{2}}+{e_{35}}{\beta^{2}}+{\omega_{0}}$ (27)
$v$ $={v_{0}}+{v_{20}}+{v_{02}}={e_{36}}{\alpha^{2}}{\eta^{4}}+\left({{e_{37}}{\alpha^{2}}+{e_{38}}{\beta^{2}}}\right){\eta^{2}}+{e_{39}}{\alpha^{2}}+{e_{310}}{\beta^{2}}+{v_{0}},$
and the coupling correction term
$\eta\Delta$ $=\eta\left({{d_{00}}+{d_{20}}{\alpha^{2}}+{d_{02}}{\beta^{2}}}\right)$ (28)
$=\eta\left[{{l_{1}}{\eta^{4}}{\alpha^{2}}+\left({{l_{2}}{\alpha^{2}}+{l_{3}}{\beta^{2}}}\right){\eta^{2}}+{l_{4}}{\alpha^{2}}+{l_{5}}{\beta^{2}}-\left({\omega_{0}^{2}-v_{0}^{2}}\right)}\right]=0$
where $a_{ij}$, $b_{ij}$, $d_{ij}$, $e_{ij}$, and $l_{i}$ are constant as provided in Appendix 5.1.
Equation (28) establishes an explicit relationship between $\eta$ and $\alpha$ and $\beta$, i.e., $\eta=\eta(\alpha,\beta)$. It obviously has a trivial solution $\eta=0$ for any values of $\alpha$ and $\beta$. In this case, the third-order solution (26) describes Lissajous orbits. With the increment of $\alpha$ and $\beta$, other non-zero real solutions will bifurcate if we have
$\Delta\left({\alpha,\beta,\eta}\right)={l_{1}}{\eta^{4}}{\alpha^{2}}+({{l_{2}}{\alpha^{2}}+{l_{3}}{\beta^{2}}}){\eta^{2}}+{l_{4}}{\alpha^{2}}+{l_{5}}{\beta^{2}}-({\omega_{0}^{2}-v_{0}^{2}})=0.$ (29)
In this case, solution (26) describes quasihalo orbits. To find the critical condition for bifurcation, we let $\eta=0$ in (29) and obtain
$\Delta\left({\alpha,\beta}\right)={l_{4}}{\alpha^{2}}+{l_{5}}{\beta^{2}}-({\omega_{0}^{2}-v_{0}^{2}})=0$ (30)
It is not difficult to calculate and verify that $l_{1}>0,l_{2}<0,l_{3}<0,l_{4}>0$ and $l_{5}<0$ for the three collinear libration points with all system parameter $\mu\in\left({0,0.5}\right]$. Hence, (30) is a hyperbolic equation where a bifurcation occurs and suitable non-zero real solutions exist for (29) when $\Delta(\alpha,\beta)$ as illustrated in Fig. 1.
Figure 1: Hyperbola for the bifurcation equation (30) with $\alpha$ and $\beta$.
Particularly, if the amplitude $\beta$ is set to zero, the third-order solution (26) describes halo orbits. It is seen from (30) that the minimum permissible value of $\alpha$ is given as (without loss of generality, only the case of $\alpha>0$ is considered in the following)
${\alpha_{\min}}=\sqrt{\left({\omega_{0}^{2}-v_{0}^{2}}\right)/{l_{4}}}.$ (31)
When $\alpha>\alpha_{\min}$ and $\beta=0$, by solving (29) we can find two solutions of $\eta$,
${\eta^{2}}=\frac{{-{l_{2}}{\alpha^{2}}-\sqrt{{l_{2}}^{2}{\alpha^{4}}-4{l_{1}}{\alpha^{2}}\left({{l_{4}}{\alpha^{2}}-\omega_{0}^{2}+v_{0}^{2}}\right)}}}{{2{l_{1}}{\alpha^{2}}}}>0$ (32)
The solution branch corresponds to northern halo orbits (Class I) for $\eta>0$ and southern halo orbits (Class II) for $\eta<0$. These results align with the classical outcomes of the third-order analytical solution of halo orbits presented in [14]. To sum up, quasihalo orbits bifurcate from Lissajous orbits when $\eta$ is a solution of (30). In particular, halo orbits bifurcate from planar Lyapunov periodic orbits when $\eta$ is a solution of (30) with $\beta$ = 0.
The third-order solution (26) serves as an initial approximation to the center manifolds of around the collinear libration points in the RTBP. A higher-order series solution is required for real space missions. In such case, the bifurcation equation is no longer a hyperbola. Figure 1 just provides a basic outline of the feasible region for $\eta$, and in fact, its real feasible region is more intricate than depicted in Fig. 1. A higher-order numerical feasible region of $\eta$ will be presented in the following section.
Figure 2: Feasible region of $\eta$ for center manifolds of $L_{1}$ in the Sun-Earth system.
Figure 3: Distribution of $\eta$ within feasible region for center manifolds of $L_{1}$.
4.2 Numerical results
The semi-analytical computation of center manifolds up to a certain order $n$ for the given system parameter $\mu$ is implemented to verify the proposed method. The analysis includes a thorough examination of the high-order bifurcation process that gives rise to quasihalo orbits from Lissajous orbits. Furthermore, the practical convergence domain of the proposed analytical solution for both Lissajous orbits and quasihalo orbits is provided. In addition, for validation purposes, the coefficients of the expansion up to the third order for center manifolds of $L_{1}$ in Sun-Earth system ($\mu$ = 3.040423398444176×$\times 10^{-6}$, $\gamma_{1}$ = 1.00109772277814×$\times 10^{-2}$) are shows in Appendix 5.2.
The third-order analytical solution in Subsection 4.1 illustrates that a feasible region of coupling coefficient $\eta$ of center manifolds is bounded by a hyperbola. Similar to the process in (30) for the third-order solution, higher-order $\Delta(\alpha,\beta)$ can be obtained during the computation of higher-order semi-analytical solutions. For each given paired amplitudes $(\alpha,\beta)$, non-trivial values of $\eta$ can be determined from $\Delta(\alpha,\beta)=0$. Then, the feasible region of $\eta$ of center manifolds around $L_{1}$ in the Sun-Earth system with different orders is presented in Fig. 2, where only the range of $\alpha\in[0,0.35]$, $\beta\in[0,0.4]$ and $\eta\in(0,3.0]$ is considered due to the symmetry of $\eta$ and the divergence of (17) for large amplitudes. The colorbar represents the number of solutions $N(\eta)$ of $\Delta(\alpha,\beta)=0$ with the range $\eta\in(0,3.0]$. It can be seen that $N$ has only two possible values $N=0$ or $N=1$, distributed on both sides of a hyperbola boundary when the order $n=3$. This result is consistent with the analytical solution in Subsection 4.1. As the order increases, parts of the region with $N=0$ and $N=1$ are replaced by the region with $N(\eta)>1$. This means that, with increasing amplitudes $\alpha$ and $\beta$, the Lissajous orbit undergoes multiple bifurcations, forming multiple families of quasihalo orbits.
Figure 4: Typical orbits around $L_{1}$ in the Sun-Earth system.
The polynomials $\Delta(\alpha,\beta)$ may encompass all the information about local bifurcation around collinear libration points in the RTBP. It is reasonable to believe that with further increases in order, the number of solutions of $\eta$ will also increase. Furthermore, it is observed that the number of solutions for $\eta$ multiplies (due to only $\eta\in(0,3.0]$ being considered, some solutions for $\eta$ are omitted here). However, larger values of $\eta$ might lead to divergence of the solution. This is because the power of $\eta$ in the coefficients of the $n$-th order terms in the series solution (17) does not exceed $2n$. To ensure the convergence of the series solution, it is required that $\eta^{2n}(\alpha^{i}\beta^{j})<1\,(i+j<n)$, which implies that $\eta$ needs to satisfy $\left|\eta\right|<\sqrt{1/{{\{\alpha,\beta\}}_{\max}}}$ at minimum, where $\{\alpha,\beta\}_{\max}$ denotes taking the maximum value between $\alpha$ and $\beta$.
Then, distributions of coupling coefficient $\eta$ values with an order $n=19$ is computed as shown in Fig. 3. The feasible region corresponds to the bottom-right part of Fig. 2. Each contour map describes the distribution of one solution of $\eta$. Empty regions indicate no solution for $\eta$. Six contour maps represent a maximum of six solutions of $\eta$ at the 19-th order. Figure 3 shows that the feasible region is largest when there is only one non-zero solution for $\eta$ (top left). Subsequently, as the number of non-zero $\eta$ solution increases, the feasible region becomes small. When the amplitude $\alpha$ and $\beta$ are small, there are no non-zero $\eta$, corresponding to the empty region in the top-left part of Fig. 3. In this case, the higher-order series (17) only describes planar and vertical Lyapunov orbits and Lissajous orbits. Figure 4 shows a typical plot of planar Lyapunov orbits by the red line ($\alpha$ = 0.05, $\beta$ = 0.0), vertical Lyapunov orbits by the black line ($\alpha$ = 0.0, $\beta$ = 0.25) and Lissajous orbits by the blue line ($\alpha$ = 0.05, $\beta$ = 0.25) in the synodic coordinate system.
Upon increasing $\alpha$ and $\beta$ to the region with one solution of $\eta$, Fig. 3 shows that $\eta$ values change from zero to non-zero but they are very small, which means the weak coupling of two degrees of freedom (in-plane and out-of-plane motions). In this case, quasihalo orbits closely resemble the corresponding Lissajous orbits, with the amplitudes identical to quasihalo orbits but $\eta$ values being zeros. This observation is evident from Fig. LABEL:fig:Lissajous_and_quasihalo_orbit, where $\eta$ = 0.08180669069. As the coupling effect ($\eta$ > 0) of the in-plane motion acts on the out-of-plane motion, a plane-symmetric Lissajous orbit (linear part in $z$-direction: $z_{1}=\beta\cos\theta_{2}$) becomes an approximately oblique-upward symmetric quasihalo orbit (linear part in $z$-direction: $z_{1}=\beta\cos\theta_{2}+\eta\alpha\cos\theta_{1}$).
With further increment in $\alpha$ and $\beta$, one $\eta$ value rapidly grows larger and more solutions for $\eta$ appear. Figure LABEL:fig:Two_quasihalo_orbits shows two quasihalo orbits with the same amplitudes. It can be seen that the quasihalo orbit with the smaller value of $\eta$ exhibits weak coupling and remains similar to Lissajous orbits. The larger value of $\eta$ leads to a significant coupling effect on the motion in the $z$-direction, resulting in a typical quasihalo orbit, as shown in Fig. LABEL:fig:Two_quasihalo_orbitsb. In this case, in-plane and out-of-plane motions are strongly coupled. As $\alpha$ and $\beta$ increase to the region shown in the bottom-left of Fig.3, four solutions for $\eta$ can be found from $\Delta=0$, i.e., $N(\eta)=4$. Figure LABEL:fig:Four_quasihalo_orbits shows four quasihalo orbits corresponding to these four solutions for the given $\alpha$ and $\beta$. It can be seen they have some strange bendings, likely due to the corresponding $\alpha$ and $\beta$ not being within the practical region of convergence.
Now, we analyze the practical convergence domain of the proposed analytical solution by comparing it with numerical solutions. Firstly, an initial condition is obtained for given amplitudes $\alpha$ and $\beta$ from the analytical solution (17). Subsequently, numerical integration of the dynamical equations is performed over a normalized time length $T=\pi$. The accuracy of the analytical solution is determined by comparing the Euclidean norm of difference in position vectors at final time between the analytical and numerically integrated solutions. Performing the same procedure for each pair of amplitudes ($\alpha$, $\beta$) within a given range yields the practical convergence domain of the solution with a specified order.
Figure LABEL:fig:Convergence_domain shows the domain of practical convergence of the proposed analytical solution up to order 35 for Lissajous orbits ($\eta=0$) and quasihalo orbits ($\eta\neq 0$) around $L_{1}$ of the Sun-Earth system, where color scale represent the normalized Euclidean norm of difference in position vectors between the analytical and numerically integrated solutions. Here Fig. LABEL:fig:Convergence_domain(a) is similar to the result in [4]. This is because, when $\eta$ is zero, no bifurcation occurs, and the analytical solution is identical to the solution of Lissajous orbits presented in [4]. However, when $\eta$ is non-zero, coupling effects between different degrees of freedom lead to the bifurcation, specifically the generation of quasihalo orbits. Figure LABEL:fig:Convergence_domain(b), for the first time, provides the actual convergence domain of the approximate analytical solution for quasihalo orbits generated by the first bifurcation from Lissajous orbits. It can be observed that the practical convergence domain for quasihalo orbits is significantly smaller than the practical convergence domain for the Lissajous orbit. This is due to the coupling in the plane direction causing a large actual amplitude in the z-direction, even when the amplitude $\beta$ is small.
Finally, let us visually illustrate the Poincaré section of the center manifold in synodic coordinate system using the solution (17). It is known that center manifolds around collinear libration points in the RTBP are four-dimensional, making direct graphical display challenging. As a byproduct of the proposed analytical method for constructing center manifolds, the dynamics inside center manifolds can be globally described in a two-dimensional Poincaré section by fixing $z=0$ with $z>0$ and the Jacobian integral $C=C_{0}$ in synodic coordinate system. First, arbitrary phase angles $\phi_{1}$ and $\phi_{2}$ and arbitrary initial time $t_{0}$ from (17) is chosen. For the chosen $C_{0}$, paired amplitudes ($\alpha$, $\beta$) and corresponding initial states ($x_{0},y_{0},z_{0}$) can be computed. Then, starting from each initial state, every points $(x_{i},y_{i})$ is plotted in the two-dimensional Poincaré section when $z_{i}=0$ within a specified time interval according to (17).
Figure LABEL:fig:Poincare_section shows two Poincaré sections projected onto the $xy$-plane of the synodic coordinate system. Each section defies a closed region. The boundary of the region is a planar Lyapunov orbit with the paired amplitudes ($\alpha_{\rm{max}}$, 0) and a fixed point on $y$-axis represents a vertical Lyapunov orbit with the paired amplitudes (0, $\beta_{\rm{max}}$), where $\alpha_{\rm{max}}$ and $\beta_{\rm{max}}$ are maximum in-plane and out-of-plane amplitudes for the selected $C_{0}$, respectively. Other circles inside this region respond to Lissajous orbits with the paired amplitudes ($\alpha$, $\beta$) where $\alpha$ < $\alpha_{\rm{max}}$ and $\beta<\beta_{\rm{max}}$. The bifurcation occurs when the increase in $C$ results in a non-zero solution $\eta\neq 0$ for the equation $\Delta=\sum\limits_{0\leq i+j\leq n}{{d_{ij}}{\alpha^{i}}{\beta^{j}}=0}$. This visual computing not only shows the emergence of well-known halo orbits from planar Lyapunov orbits but also illustrates the generation of quasi-periodic orbits from Lissajous orbits.
5 Conclusions
Understanding the coupling between in-plane and out-of-plane motions and its connection to the bifurcation of halo/quasihalo orbits is crucial for unraveling the mechanism governing their generation and obtaining their unified analytical solution. This paper introduced a novel concept of coupling coefficients into the RTBP for the first time, incorporated a new correction term into the RTBP equation to characterize this coupling effect and successfully derived a unified analytical solution for the center manifolds of collinear libration points in the RTBP. When the bifurcation equation $\Delta=0$ has no real solutions, the proposed series solution describes planar Lyapunov orbits, vertical Lyapunov orbits and Lissajous orbits. When the equation $\Delta=0$ has only zero solutions, the bifurcation occurs. When the equation $\Delta=0$ yields non-zero real solution, the proposed series solution further describes both halo orbits and quasihalo orbits, with the latter being constructed analytically for the first time. Remarkably, larger amplitudes lead to multiple solutions of the equation $\Delta=0$ with the higher-order series solution, indicating the discovery of multiple bifurcations of quasihalo orbits. These solutions can all be explicitly calculated, although some may fall outside the domain of practical convergence due to the order limitation. Finally, it is noted that the coupling efficient $\eta$ and the equation $\Delta=0$ serve as conditions for bifurcation. Thus, this method can also be referred to as an analytical bifurcation calculation method.
Appendix
5.1 Coefficients of third-order analytical solution
${a_{21}}=\frac{{3{c_{3}}({\kappa^{2}}-2)}}{{4(1+2{c_{2}})}},{a_{22}}=\frac{{3{c_{3}}}}{{4(1+2{c_{2}})}},{a_{23}}=-\frac{{3{c_{3}}}}{4}\frac{{({\kappa^{2}}+2)(\varpi_{20}^{2}-{c_{2}}+1)+4\kappa{\varpi_{20}}}}{{{s_{20}}}}$
${a_{24}}=\frac{{3{c_{3}}}}{4}\frac{{(\varpi_{20}^{2}-{c_{2}}+1)}}{{{s_{20}}}},{a_{25}}=\frac{{3{c_{3}}}}{2}\frac{{\varpi_{11}^{2}-{c_{2}}+1}}{{{s_{11}}}},{a_{26}}=\frac{{3{c_{3}}}}{2}\frac{{\varpi_{1-1}^{2}-{c_{2}}+1}}{{{s_{1-1}}}}$
${a_{27}}=\frac{{3{c_{3}}}}{{4(1+2{c_{2}})}},{a_{28}}=\frac{{3{c_{3}}}}{4}\frac{{\varpi_{02}^{2}-{c_{2}}+1}}{{{s_{02}}}}$
$\ {a_{31}}=\frac{{3{c_{3}}\left({\varpi_{30}^{2}-{c_{2}}+1}\right){d_{24}}}}{{2{s_{30}}}},{a_{32}}=-\frac{{\left({\varpi_{30}^{2}-{c_{2}}+1}\right){t_{1}}-2{\varpi_{30}}{t_{2}}}}{{{s_{30}}}}$
$\ {a_{33}}=-\frac{{\left({\varpi_{30}^{2}-{c_{2}}+1}\right){t_{3}}-3{\varpi_{30}}{t_{4}}}}{{{s_{30}}}},{a_{34}}=-\frac{{\left({v_{00}^{2}-{c_{2}}+1}\right){t_{5}}}}{{{s_{01}}}}$
$\ {a_{35}}=-\frac{{\left({v_{00}^{2}-{c_{2}}+1}\right){t_{6}}-2{v_{00}}{t_{7}}}}{{{s_{01}}}},{a_{36}}{\rm{=}}-\frac{{\left({\varpi_{21}^{2}-{c_{2}}+1}\right){t_{8}}}}{{{s_{21}}}}$
$\ {a_{37}}{\rm{=}}-\frac{{\left({\varpi_{21}^{2}-{c_{2}}+1}\right){t_{9}}-2{\varpi_{21}}{t_{10}}}}{{{s_{21}}}},{a_{38}}=-\frac{{\left({\varpi_{2-1}^{2}-{c_{2}}+1}\right){t_{11}}}}{{{s_{2-1}}}}$
$\ {a_{39}}=-\frac{{\left({\varpi_{2-1}^{2}-{c_{2}}+1}\right){t_{12}}-2{\varpi_{2-1}}{t_{13}}}}{{{s_{2-1}}}},{a_{310}}=-\frac{{\left({\varpi_{12}^{2}-{c_{2}}+1}\right){t_{14}}}}{{{s_{12}}}}$
$\ {a_{311}}=\frac{{\left({\varpi_{12}^{2}-{c_{2}}+1}\right){t_{15}}-2{\varpi_{12}}{t_{16}}}}{{{s_{12}}}},{a_{312}}=-\frac{{\left({\varpi_{1-2}^{2}-{c_{2}}+1}\right){t_{17}}}}{{{s_{1-2}}}}$
$\ {a_{313}}=\frac{{\left({\varpi_{1-2}^{2}-{c_{2}}+1}\right){t_{18}}-2{\varpi_{1-2}}{t_{19}}}}{{{s_{1-2}}}},{a_{314}}=\frac{{3{c_{3}}\left({\varpi_{01}^{2}-{c_{2}}+1}\right)\left({2{d_{29}}+{d_{210}}}\right)}}{{2{s_{01}}}}$
$\ {a_{315}}=\frac{{3{c_{3}}\left({\varpi_{01}^{2}-{c_{2}}+1}\right){d_{210}}}}{{2{s_{03}}}}$
$\ {b_{21}}=\frac{{3{c_{3}}}}{2}\frac{{\kappa\varpi_{20}^{2}+({\kappa^{2}}+2){\varpi_{20}}+\kappa(2{c_{2}}+1)}}{{{s_{20}}}},{b_{22}}=-\frac{{3{c_{3}}}}{2}\frac{{{\varpi_{20}}}}{{{s_{20}}}}$
$\ {b_{23}}=\frac{{-3{c_{3}}{\varpi_{11}}}}{{{s_{11}}}},{b_{24}}=\frac{{-3{c_{3}}{\varpi_{1-1}}}}{{{s_{1-1}}}},{b_{25}}=-\frac{{3{c_{3}}}}{2}\frac{{{\varpi_{02}}}}{{{s_{02}}}}$
${b_{31}}=-\frac{{3{c_{3}}(\kappa{\omega_{0}}+1)(2{d_{22}}+{d_{24}})}}{{2{s_{1}}}},{b_{32}}=\frac{{(\kappa{\omega_{0}}+1){t_{20}}-({\omega_{0}}+\kappa){t_{21}}}}{{{s_{1}}}}$
${b_{33}}=\frac{{(\kappa{\omega_{0}}+1){t_{22}}-({\omega_{0}}+\kappa){t_{23}}}}{{{s_{1}}}},{b_{34}}=-\frac{{6{c_{3}}{\varpi_{30}}{d_{24}}}}{{2{s_{30}}}}$
${b_{35}}=\frac{{2{\varpi_{30}}{t_{1}}-\left({\varpi_{30}^{2}+2{c_{2}}+1}\right){t_{2}}}}{{{s_{30}}}},{b_{36}}=\frac{{2{\varpi_{30}}{t_{3}}-\left({\varpi_{30}^{2}+2{c_{2}}+1}\right){t_{4}}}}{{{s_{30}}}}$
${b_{37}}=\frac{{2{v_{00}}{t_{5}}}}{{{s_{10}}}},{b_{38}}=\frac{{2{v_{00}}{t_{6}}-\left({v_{00}^{2}+2{c_{2}}+1}\right){t_{7}}}}{{{s_{01}}}}$
${b_{39}}=\frac{{2{\varpi_{21}}{t_{8}}}}{{{s_{21}}}},{b_{310}}=\frac{{2{\varpi_{21}}{t_{9}}-\left({\varpi_{21}^{2}+2{c_{2}}+1}\right){t_{10}}}}{{{s_{21}}}}$
${b_{311}}=\frac{{2{\varpi_{2-1}}{t_{11}}}}{{{s_{2-1}}}},{b_{312}}=\frac{{2{\varpi_{2-1}}{t_{12}}-\left({\varpi_{2-1}^{2}+2{c_{2}}+1}\right){t_{13}}}}{{{s_{2-1}}}}$
${b_{313}}=\frac{{(\kappa{\omega_{0}}+1){t_{24}}}}{{{s_{1}}}},{b_{314}}=\frac{{(\kappa{\omega_{0}}+1){t_{25}}-({\omega_{0}}+\kappa){t_{26}}}}{{{s_{1}}}}$
${b_{315}}=\frac{{2{\varpi_{12}}{t_{14}}}}{{{s_{12}}}},{b_{316}}=\frac{{2{\varpi_{12}}{t_{15}}-\left({\varpi_{12}^{2}+2{c_{2}}+1}\right){t_{16}}}}{{{s_{12}}}}$
${b_{317}}=\frac{{2{\varpi_{1-2}}{t_{17}}}}{{{s_{1-2}}}},{b_{318}}=\frac{{2{\varpi_{1-2}}{t_{18}}-\left({\varpi_{1-2}^{2}+2{c_{2}}+1}\right){t_{19}}}}{{{s_{1-2}}}}$
${b_{319}}=-\frac{{3{c_{3}}{v_{00}}(2{d_{29}}+{d_{210}})}}{{{s_{01}}}},{b_{320}}=-\frac{{3{c_{3}}{d_{210}}{\varpi_{03}}}}{{{s_{03}}}}$
${d_{21}}=-\frac{{3{c_{3}}}}{{2{c_{2}}}}-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}}}{a_{21}},{d_{22}}=-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}}}{a_{22}}$
${d_{23}}=-\frac{{3{c_{3}}}}{{2({c_{2}}-\varpi_{20}^{2})}}-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}-\varpi_{20}^{2}}}{a_{23}},{d_{24}}=-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}-\varpi_{20}^{2}}}{a_{24}}$
${d_{25}}=-\frac{{3{c_{3}}}}{{2({c_{2}}-\varpi_{11}^{2})}},{d_{26}}=-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}-\varpi_{11}^{2}}}{a_{25}}$
${d_{27}}=-\frac{{3{c_{3}}}}{{2({c_{2}}-\varpi_{1-1}^{2})}},{d_{28}}=-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}-\varpi_{1-1}^{2}}}{a_{26}}$
${d_{29}}=-\frac{{3{c_{3}}(\omega_{0}^{2}-v_{0}^{2})}}{{4(1+2{c_{2}}){c_{2}}}},{d_{210}}=-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}-\varpi_{02}^{2}}}{a_{28}}$
${d_{31}}=-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}-\varpi_{30}^{2}}}{a_{31}},{d_{32}}=\frac{{{t_{27}}}}{{{c_{2}}-\varpi_{30}^{2}}}-\frac{{\left({\omega_{0}^{2}-v_{0}^{2}}\right){a_{32}}}}{{{c_{2}}-\varpi_{30}^{2}}}$
${d_{33}}=\frac{{{t_{28}}}}{{{c_{2}}-\varpi_{30}^{2}}}-\frac{{\left({\omega_{0}^{2}-v_{0}^{2}}\right){a_{33}}}}{{{c_{2}}-\varpi_{30}^{2}}},{d_{34}}=-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}-\varpi_{21}^{2}}}{a_{36}}$
${d_{35}}=\frac{{{t_{29}}-\left({\omega_{0}^{2}-v_{0}^{2}}\right){a_{37}}}}{{{c_{2}}-\varpi_{21}^{2}}},{d_{36}}=\frac{{{t_{30}}}}{{{c_{2}}-\varpi_{21}^{2}}},{d_{37}}=-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}-\varpi_{2-1}^{2}}}{a_{38}}$
${d_{38}}=\frac{{{t_{31}}}}{{{c_{2}}-\varpi_{2-1}^{2}}}-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}-\varpi_{2-1}^{2}}}{a_{39}},{d_{39}}=\frac{{{t_{32}}}}{{{c_{2}}-\varpi_{2-1}^{2}}}$
${d_{310}}=-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}-\varpi_{12}^{2}}}{a_{310}},{d_{311}}=\frac{{{t_{33}}}}{{{c_{2}}-\varpi_{12}^{2}}}-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}-\varpi_{12}^{2}}}{a_{311}}$
${d_{312}}=-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}-\varpi_{1-2}^{2}}}{a_{312}},{d_{313}}=\frac{{{t_{34}}}}{{{c_{2}}-\varpi_{1-2}^{2}}}-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}-\varpi_{1-2}^{2}}}{a_{313}}$
${d_{314}}=-\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{{c_{2}}-\varpi_{03}^{2}}}{a_{315}},{d_{315}}=\frac{3}{8}\frac{{4{c_{3}}{a_{28}}+{c_{4}}}}{{{c_{2}}-\varpi_{03}^{2}}}$
${e_{31}}=\frac{1}{2}\frac{{3{c_{3}}(\omega_{0}^{2}-{c_{2}}+1)(2{d_{22}}+{d_{24}})}}{{2{s_{1}}}},{e_{32}}=-\frac{1}{2}\frac{{(\omega_{0}^{2}-{c_{2}}+1){t_{20}}-2{\omega_{0}}{t_{21}}}}{{{s_{1}}}}$
${e_{33}}=-\frac{1}{2}\frac{{(\omega_{0}^{2}-{c_{2}}+1){t_{22}}-2{\omega_{0}}{t_{23}}}}{{{s_{1}}}},{e_{34}}=\frac{{\left({\omega_{0}^{2}-v_{0}^{2}}\right){a_{34}}}}{{2{v_{00}}}}$
${e_{35}}=-\frac{{{t_{38}}-\left({\omega_{0}^{2}-v_{0}^{2}}\right){a_{35}}}}{{2{v_{00}}}},{e_{36}}=\frac{{\frac{3}{2}{c_{3}}\left({2{a_{21}}+{d_{25}}+{d_{27}}}\right)+\frac{3}{4}\left({4-{\kappa^{2}}}\right){c_{4}}}}{{2{v_{00}}}}$
${e_{37}}=-\frac{1}{2}\frac{{(\omega_{0}^{2}-{c_{2}}+1){t_{24}}}}{{{s_{1}}}},{e_{38}}=-\frac{1}{2}\frac{{(\omega_{0}^{2}-{c_{2}}+1){t_{25}}-{\omega_{0}}{t_{26}}}}{{{s_{1}}}}$
${e_{39}}=\frac{{\omega_{0}^{2}-v_{0}^{2}}}{{2{v_{00}}}}{a_{314}},{e_{310}}=\frac{{\frac{3}{2}{c_{3}}(2{a_{27}}+{a_{28}})-\frac{9}{8}{c_{4}}}}{{2{v_{00}}}}$
${l_{1}}=-2{\omega_{0}}{e_{31}},{l_{2}}=-2{\omega_{0}}{e_{32}}+{t_{35}},{l_{3}}=-2{\omega_{0}}{e_{34}}-{t_{36}}$
${l_{4}}=-2{\omega_{0}}{e_{33}},{l_{5}}=-2{\omega_{0}}{e_{35}}-{t_{37}}$
where
${\varpi_{km}}=k{\omega_{0}}+m{v_{0}}$
${s_{km}}=\varpi_{km}^{2}(\varpi_{km}^{2}-2+{c_{2}})-(2{c_{2}}+1)({c_{2}}-1)$
${s_{1}}=\omega_{0}^{3}-\kappa{\omega_{0}}^{2}-({c_{2}}+1){\omega_{0}}-\kappa({c_{2}}-1),$
${t_{1}}=\frac{3}{2}{c_{3}}\left({2{a_{24}}+\kappa{b_{22}}-{d_{23}}}\right)-\frac{3}{2}{c_{4}}$
${t_{2}}=-\frac{3}{2}{c_{3}}\left({\kappa{a_{24}}+{b_{22}}}\right)+\frac{3}{8}{c_{4}}\kappa$
${t_{3}}=\frac{3}{2}{c_{3}}\left({2{a_{23}}+\kappa{b_{21}}}\right)+\frac{1}{2}{c_{4}}({2+3{\kappa^{2}}})$
${t_{4}}=-\frac{3}{2}{c_{3}}\left({\kappa{a_{23}}+{b_{21}}}\right)-\frac{3}{8}{c_{4}}\kappa,{t_{5}}=-\frac{3}{2}{c_{3}}\left({2{d_{22}}+{d_{26}}+{d_{28}}}\right)$
${t_{6}}=\frac{3}{2}{c_{3}}\left({2{a_{25}}+2{a_{26}}-\kappa{b_{23}}-\kappa{b_{24}}-2{d_{21}}-{d_{25}}-{d_{27}}}\right)-6{c_{4}}$
${t_{7}}=\frac{3}{2}{c_{3}}\left({\kappa{a_{25}}-\kappa{a_{26}}-{b_{23}}+{b_{24}}}\right)$
${t_{8}}=-\frac{3}{2}{c_{3}}\left({{d_{24}}+{d_{26}}}\right)$
${t_{9}}=\frac{3}{2}{c_{3}}\left({2{a_{25}}+\kappa{b_{23}}-{d_{23}}-{d_{25}}}\right)-3{c_{4}}$
${t_{10}}=-\frac{3}{2}{c_{3}}\left({\kappa{a_{25}}+{b_{23}}}\right)+\frac{3}{4}\kappa{c_{4}}$
${t_{11}}=-\frac{3}{2}{c_{3}}\left({{d_{24}}+{d_{28}}}\right)$
${t_{12}}=-\frac{3}{2}{c_{3}}\left({{d_{23}}+{d_{27}}}\right)-3{c_{4}}$
${t_{13}}=-\frac{3}{2}{c_{3}}\left({\kappa{a_{26}}+{b_{24}}}\right)-\frac{3}{4}\kappa{c_{4}}$
${t_{14}}=-\frac{3}{2}{c_{3}}\left({{d_{26}}+{d_{30}}}\right)$
${t_{15}}=\frac{3}{2}{c_{3}}\left({2{a_{28}}+\kappa{b_{25}}-{d_{25}}}\right)-\frac{3}{2}{c_{4}}$
${t_{16}}=-\frac{3}{2}{c_{3}}\left({\kappa{a_{28}}+{b_{25}}}\right)+\frac{3}{8}{c_{4}}\kappa,{t_{17}}=-\frac{3}{2}{c_{3}}\left({{d_{28}}+{d_{30}}}\right)$
${t_{18}}=\frac{3}{2}{c_{3}}\left({2{a_{28}}-\kappa{b_{25}}-{d_{27}}}\right)-\frac{3}{2}{c_{4}}$
${t_{19}}=-\frac{3}{2}{c_{3}}\left({\kappa{a_{28}}-{b_{25}}}\right)+\frac{3}{8}{c_{4}}\kappa$
${t_{20}}=\frac{3}{2}{c_{3}}\left({4{a_{22}}+2{a_{24}}-\kappa{b_{22}}-2{d_{21}}-{d_{23}}}\right)-\frac{9}{2}{c_{4}}$
${t_{21}}=-\frac{3}{2}{c_{3}}\left({2\kappa{a_{22}}-\kappa{a_{24}}+{b_{22}}}\right)+\frac{3}{8}{c_{4}}\kappa$
${t_{22}}=\frac{3}{2}{c_{3}}\left({4{a_{21}}+2{a_{23}}-\kappa{b_{21}}}\right)+\frac{3}{2}{c_{4}}({2-{\kappa^{2}}})$
${t_{23}}=-\frac{3}{2}{c_{3}}\left({2\kappa{a_{21}}-\kappa{a_{23}}+{b_{21}}}\right)-\frac{3}{8}{c_{4}}\kappa({4-3{\kappa^{2}}})$
${t_{24}}=-\frac{3}{2}{c_{3}}\left({{d_{26}}+{d_{28}}+2{d_{29}}}\right)$
${t_{25}}=\frac{3}{2}{c_{3}}\left({4{a_{27}}-{d_{25}}-{d_{27}}}\right)-3{c_{4}}$
${t_{26}}=-\frac{3}{4}\kappa\left({4{c_{3}}{a_{27}}-{c_{4}}}\right)$
${t_{27}}=-\frac{3}{2}{c_{3}}({a_{24}}+{d_{24}})+\frac{3}{8}{c_{4}}$
${t_{28}}=-\frac{3}{2}{c_{3}}\left({{a_{23}}+{d_{23}}}\right)-\frac{3}{8}{c_{4}}({4+{\kappa^{2}}})$
${t_{29}}=-\frac{3}{2}{c_{3}}\left({{a_{24}}+{a_{25}}+{d_{26}}}\right)+\frac{9}{8}{c_{4}}$
${t_{30}}=-\frac{3}{2}{c_{3}}\left({{a_{23}}+{d_{25}}}\right)-\frac{3}{8}{c_{4}}({4+{\kappa^{2}}})$
${t_{31}}=-\frac{3}{2}{c_{3}}\left({{a_{24}}+{a_{26}}+{d_{28}}}\right)+\frac{9}{8}{c_{4}}$
${t_{32}}=-\frac{3}{2}{c_{3}}\left({{a_{23}}+{d_{27}}}\right)-\frac{3}{8}{c_{4}}({4+{\kappa^{2}}})$
${t_{33}}=-\frac{3}{2}{c_{3}}\left({{a_{25}}+{a_{28}}+{d_{210}}}\right)+\frac{9}{8}{c_{4}}$
${t_{34}}=-\frac{3}{2}{c_{3}}\left({{a_{26}}+{a_{28}}+{d_{210}}}\right)+\frac{9}{8}{c_{4}}$
${t_{35}}=-\frac{3}{2}{c_{3}}\left({2{a_{22}}+{a_{24}}+2{d_{22}}+{d_{24}}}\right)+\frac{9}{8}{c_{4}}$
${t_{36}}=-\frac{3}{2}{c_{3}}\left({2{a_{21}}+{a_{23}}+2{d_{21}}+{d_{23}}}\right)-\frac{3}{8}{c_{4}}({12-{\kappa^{2}}})$
${t_{37}}=-\frac{3}{2}{c_{3}}\left({{a_{25}}+{a_{26}}+2{a_{27}}+2{d_{29}}}\right)+\frac{9}{4}{c_{4}}$
${t_{38}}=-\frac{3}{2}{c_{3}}\left({2{a_{22}}+{a_{25}}+{a_{26}}+{d_{26}}+{d_{28}}}\right)+\frac{9}{4}{c_{4}}.$
5.2 Coefficient of expansion for center manifolds of $L_{1}$ in Sun-Earth system (up to order 3)
$\boldsymbol{i}$ $\boldsymbol{j}$ $\omega_{ij}$ $v_{ij}$ $d_{ij}$
0 0 2.0864535642231 2.01521066299663 -0.292214459403954
[…]
$\boldsymbol{i}$ $\boldsymbol{j}$ $\boldsymbol{k}$ $\boldsymbol{m}$ $x_{ijkm}$ $y_{ijkm}$ $z_{ijkm}$
1 0 1 0 1 -3.22926825193629 $\eta$
0 1 0 1 1
[…]